OpenStudy (anonymous):

if you need to evaluate the integral but its in fraction form... how can one do this ? Problem: ∫ xdx / (7x^2 + 3)^5

OpenStudy (anonymous):

Have you tried a u substitution?

OpenStudy (anonymous):

Try, u = 7x^2+3 and see what you get.

OpenStudy (anonymous):

no i haven't tried u substition

OpenStudy (anonymous):

okay so u'/u right?

OpenStudy (anonymous):

u = 7x^2+3 du=14xdx du/14 = x dx

OpenStudy (anonymous):

okay that looks like what i got, except how did you get du/14 =xdx ?

OpenStudy (anonymous):

du/dx = 14x du/14 = x dx

OpenStudy (anonymous):

i mean why do you need to do that?

OpenStudy (anonymous):

1/14 is just a constant, you can take this out and x dx is in your numerator you want du to equal that.

OpenStudy (anonymous):

i see so i will plug that in to my numerator..what about the x besides the dx in the numerator ?

OpenStudy (anonymous):

du/14 = xdx So in your numerator xdx becomes du/14 and now you'll be integrating respect to du.

OpenStudy (anonymous):

okay so I will integrate as I normally do ? can you show me one part to this process so i can get an idea?

OpenStudy (anonymous):

du/14/(7x^2 + 3)^5 ?

OpenStudy (anonymous):

\[\int\limits \frac{ x }{ (7x^2+3)^5 }dx~~~u=7x^2+3\implies du = 14xdx \implies \frac{ du }{ 14 }=xdx\]

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ (u)^5 } \frac{ du }{ 14 } \implies \frac{ 1 }{ 14 } \int\limits \frac{ du }{ u^5 }\]

OpenStudy (anonymous):

Now integrate 1/u^5

OpenStudy (anonymous):

hmmm okay! 1/u^5 is 1/u^6?

OpenStudy (anonymous):

mhm?

OpenStudy (anonymous):

or is it u^6/6

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ u^5 } du \implies \int\limits u^{-5} du \implies \frac{ u^{-5+1} }{ -5+1 }+C\]

OpenStudy (anonymous):

Now plug u back in

OpenStudy (anonymous):

okay okay, completely forgot about it being in the denominator means a negative exponent

OpenStudy (anonymous):

No worries :)

OpenStudy (anonymous):

plug it back into what now? I'm sorry i'm slow..:(

OpenStudy (anonymous):

the equation!

OpenStudy (anonymous):

\[\frac{ 1 }{ 14 }\frac{ u^{-4} }{ -4 }+C \implies \frac{ 1 }{ 14 }\frac{ 1 }{ -4(u)^4 } \implies \frac{ 1 }{ 14 }\frac{ 1 }{ -4(7x^2+3) }+C\] forgot to bring the 1/14 earlier :P bring that with you

OpenStudy (anonymous):

So your final answer is \[-\frac{ 1 }{ 56(7x^2+3 )}+C\]

OpenStudy (anonymous):

Did that make sense?

OpenStudy (anonymous):

Oh okay thank you so mcuh @iambatman

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

honestly i need more practice with u-substitution but yes you made much more easier !

OpenStudy (anonymous):

It takes a while to get used to, but no problem! :)

OpenStudy (anonymous):

ok good i did have one last question

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

ok when you plug the du back into the equation, i noticed the denominator part of it sort of "disappeared" ?

OpenStudy (anonymous):

It didn't disappear we just made a substitution for all the stuff with x in it.